# Discovering modular solutions that generalize compositionally

### v1

@@ A.1.2 Irreducibility condition for ReLU nonlinearity
  • There exists no output weight $a_{k}$ that is 0
@@
  • If a pair of negatively colinear input weight exists, by denoting $k,l$ the corresponding hidden neuron indices, there are no subset $\mathcal{I}$ of $1..h]\setminus\{k,l\}$such that $a_{k}{\bm{w}}_{k}=-\sum_{i\in\mathcal{I}}a_{i}{\bm{w}}_{i}$ or $a_{l}{\bm{w}}_{l}=-\sum_{i\in\mathcal{I}}a_{i}{\bm{w}}_{i}$
@@ A.2 Multi-task parameter identification
We now come back to the multi-task setting, where the teacher generates a task-specific network using a hypernetwork conditioned on a task latent variable. Specifically, the weights of the second layer are fixed across the tasks, but the weights of the first layer are obtained as a linear combination of the modules: $\hat{{\bm{W}}}(\hat{\Theta},{\bm{z}})=\sum_{m}^{M}z^{(m)}\hat{\Theta}^{(m)}$. See section 3.1 for the full description.
@@
A.2.1 Multi-task parameter identification
We provide results corresponding to ReLU MLP as well as MLPs where the activation function belongs to a general class of smooth functions, for student networks that can be wider than the teacher network. Theorem 2 in the main text is a special case of the following results.
ReLU MLP
@@ A.2.1 Multi-task parameter identification
Assume $\mathcal{P}_{x}$ has full support in the input space, $\leq n, M^=M\hat${M}=M, $\hat{h}\geq h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one $\bm{z}$} such that no two rows of $W(\Theta,{\bm{z}})$ are colinear. Finally, assume $\psi$ is the ReLU nonlinearity. Then, $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ for almost any $\bm{z}$} is equivalent to the existence of an invertible matrix $\bm{F}$}, a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ of $1..\hat{h}$] such that we have
@@
Assume there exists one point ${\bm{z}}_{0}$ such that $W(\Theta,{\bm{z}}_{0})$ has no two rows that are colinear. Then, for any linear subspace $U$ containing ${\bm{z}}_{0}$, $W(\Theta,{\bm{z}})$ has no two rows that are colinear for ${\bm{z}}\in U$ almost everywhere [^2] By ${\bm{z}}\in U$ almost everywhere” we imply using the Lebesgue measure. In particular, the set of points $\bm{z}$} in $U$ such that $W(\Theta,{\bm{z}})$ has at least two colinear rows form a finite union of strict linear subspace of $U$.
@@
Recall that the $i$th row of $W(\Theta,z_{0})$ is given by $\Theta_{i}z_{0}$. Consider a pair $i,j$ with $i\neq j$. As no two rows in $W(\Theta,z_{0})$ are colinear, there exists only a finite number of $\alpha\in\mathbb{R}$ such that the equation $(\Theta_{i}-\alpha\Theta_{j}){\bm{z}}=0$ yields solutions of $\bm{z}$} that are not in $\text{Ker }\Theta_{i}\cap\text{Ker }\Theta_{j}$. Given each of such $\alpha$, the set of $\bm{z}$} that verifies the equation form a linear subspace. Taking everything together, the set of $\bm{z}$} such that $W(\Theta,{\bm{z}})$ has at least two colinear elements is a finite union of linear subspaces. Due to the existence of ${\bm{z}}_{0}$ we know that these linear subspaces must not include any linear subspace containing ${\bm{z}}_{0}$ and the claim follows. ∎
@@
By assumption, there exists at least one point ${\bm{z}}_{0}$ such that $W(\Theta,{\bm{z}}_{0})$ has no two rows that are colinear. Following Lemma 6, for almost every $\bm{z}$} in $\mathbb{R}^{M}$, the first layer weights have no two rows that are colinear. For all $\bm{z}$}, we fix $\hat{\bm{z}}\in\arg\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})$. By assumption, $L(\hat{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ for almost every $\bm{z}$}. Considering the intersection of such $\bm{z}$} and those that generate first layer weights with no two colinear rows, we can apply Theorem 4 for almost every $\bm{z}$}. For each such $\bm{z}$} the theorem guarantees us a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ such that equation 15-equation 22 hold for the generated weights.
@@
Because there exists a full rank matrix in the linear span of the $(\Theta_{i})_{1\leq i\leq h}$, then clearly we can find some scalar coefficients $(\lambda_{i})_{i\in[1..h]}$ and linearly combine the equations 46 over $i$ such that we obtain an equation of the form $^{\bm{A}}{\bm{z}}={\bm{B}}\hat{\bm{z}}$ for all ${\bm{z}}\in\mathbb{R}^{M}$ and corresponding $\hat{\bm{z}}$, where $\bm{A}$} is full rank. Necessarily then, $\bm{B}$} is full rank, and multiplying by its pseudoinverse yields $^{\bm{F}}{\bm{z}}=\hat{\bm{z}}$ where $\bm{F}$} is an invertible matrix. $\hat{\bm{z}}$ for which the equations hold is thus a linear function of $\bm{z}$}, which also implies the constancy of the various parameters $\alpha$ w.r.t. $\bm{z}$}.
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $\leq n, M^=M\hat${M}=M, $\hat{h}=h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one $\bm{z}$} such that no two rows of $W(\Theta,{\bm{z}})$ are colinear. Finally, assume $\psi$ is the ReLU nonlinearity. Then, $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ for almost any $\bm{z}$} is equivalent to the existence of an invertible matrix $\bm{F}$}, a partition ${\cal S}=({\cal S}_{1},{\cal S}_{2})$ and a permutation $\sigma$ of $1..\hat{h}$] such that we have
@@
MLP with infinitely differentiable activation function
@@
[…]
@@
    • no output weight $a_{k}$ is 0
    • no input weight ${\bm{w}}_{k}$ is 0
@@
Then there exists a partition $(\mathcal{S}_{i})_{1\leq i\leq h}\oplus(\mathcal{S}^{\prime}_{i})_{1\leq i\leq h^{\prime}}$ of $[1..h]$ for some $h^{\prime}$ such that we have
@@
In essence, if the teacher parameters are irreducible, we learn the same parameters up to permutation when the hidden layer is of the same width. If we allow the student to have a wider width, the over-parameterization can either:
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $\leq n, M^=M\hat${M}=M, $\hat{h}\geq h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one $\bm{z}$} such that no two rows of $W(\Theta,{\bm{z}})$ are identical. Finally, assume the activation function $\psi$ satisfies $\psi\in\mathcal{C}^{\infty}$, and $\psi^{(n)}(0)\neq 0$ for infinitely many even and odd values of $n$. Then, $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ for all $\bm{z}$} is equivalent to the existence of an invertible matrix $\bm{F}$} and a partition $(\mathcal{S}_{i})_{1\leq i\leq h}\oplus(\mathcal{S}^{\prime}_{i})_{1\leq i\leq h^{\prime}}$ for some $h^{\prime}$ of $1,...\hat{h}$] such that
@@
A.2.2 Failure examples of compositional generalization in multi-task teacher-student
@@
We wish to uncover in the following the conditions under which a student might achieve composition generalization when trained on a training set of tasks. This subsection provides some failure cases where perfectly learning the training tasks does not transfer to unseen tasks. It does not provide proofs, but rather an intuition on the later stated theorems and the assumptions they require.
In what follows, $(e_{m})$ will be the canonical basis of $\mathbb{R}^{M}$.
@@
[…]
@@
We can consider a student with two neurons and two modules. Let $\Theta_{1}=(A|B)$ and $\Theta_{2}=(C|D)$ for four linearly independent vectors $A,B,C,D$, and output weights all equal $a_{1}=a_{2}=\lambda$. If we only consider tasks defined by masks $\{e_{1},e_{2}\}$, a valid student is given by $\hat{\Theta}_{1}=(A|D)$ and $\hat{\Theta}_{2}=(C|B)$ and the same output weights. This student does however not generalize to the task described by mask $e_{1}+e_{2}$
@@ A.2.3 Compositional generalization
We are now ready to state and prove the theorem. The proof for the general class of smooth activation function used in Appendix A.2.1 follows the same steps, we omit the details.
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $\leq n, M^=M\hat${M}=M, $\hat{h}=h$, that $\sum_{i}a_{i}\Theta_{i}$ is full rank, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that for all modules $\Theta^{(m)}$, no two rows are colinear, that we can extract a family $Z_{k}$ from the support $\mathcal{Z}$ of $\mathcal{P}_{z}$ with the above property, and that $\mathcal{P}_{z}$ has connected and compositional support. Then,
$\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})\right]=0\implies\forall{\bm{z}},\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ (61)
@@
[…]
@@
In a first step we prove the existence of an invertible matrix $\bm{F}$}, and for each $k$, the existence of permutations $\sigma_{k}$ of $[1..h]$, and partitions $\mathcal{S}_{1}^{k},\mathcal{S}_{2}^{k}$ of $[1..h]$ such that $\forall{\bm{z}}\in\operatorname{span}(Z_{k})$,
@@
We start by arguing that we may assume without loss of generality that all $z\in Z_{k}$ satisfy some nice properties. First we claim that we may assume that all points $z\in Z_{k}$ satisfy $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$. This holds as $\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})\right]=0$ and our assumption that $\mathbb{P}_{{\bm{z}}\sim\mathcal{P}_{z}}[{\bm{z}}\in Z_{k}]>0$ implies that the set of points $\bm{z}$} of $Z_{k}$ such that $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})\neq 0$ has zero probability.
@@
We remind the reader that if a surjective function $f:A\rightarrow B$ between two metric spaces $A,B$ equipped by metrics $\mu,\mu^{\prime}$ verifies that all pre-image of nullsets are nullsets, i.e. $\forall X\subset B:\mu^{\prime}(X)=0\implies\mu(f^{-1}(X))=0$, then we have that if some property is true for almost every ${\bm{x}}\in B$, it must be true for $f({\bm{x}})$ for almost every ${\bm{x}}\in f^{-1}(B)=A$. If the above 2 points are true, then by application of this reasoning, we get that for almost any $({\bm{z}}_{e})_{e}\in S^{d+1}\setminus C$, there exists $(\lambda_{e})_{e}$ such that property 74 holds for $(\lambda_{e}{\bm{z}}_{e})_{e\in[1..d+1]}$. Since it can be shown that the measure of $S^{d+1}\setminus C$ is 0, we have that the assertion holds almost everywhere on $S^{d+1}$, which would conclude the proof.
@@
[…]
@@
Fix some set $Z_{k}$. As in the continuous case we may assume without loss of generality that all points $\bm{z}$} in $Z_{k}$ satisfy that no two rows of ${\bm{W}}(\Theta,{\bm{z}})$ are colinear and that they achieve zero loss, i.e. $\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$.
@@
For all $\bm{z}$}, we fix $\hat{\bm{z}}\in\arg\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})$. As argued above already, it follows from Theorem 3 that for any $\bm{z}$} in $\mathcal{F}_{k}$, there exists a permutation and partition of $[1..h]$, $\sigma_{{\bm{z}}}$ and $({\cal S}_{1}({\bm{z}}),{\cal S}_{2}({\bm{z}}))$ such that we have

### v2

@@ A.1.2 Irreducibility condition for ReLU nonlinearity
  • There exists no output weight $a_{k}$ that is 00
@@
  • If a pair of negatively colinear input weight exists, by denoting $k,l$ the corresponding hidden neuron indices, there are no subset $\mathcal{I}$ of $[1..h]\setminus\{k,l\}$such that $a_{k}{\bm{w}}_{k}=-\sum_{i\in\mathcal{I}}a_{i}{\bm{w}}_{i}$ or $a_{l}{\bm{w}}_{l}=-\sum_{i\in\mathcal{I}}a_{i}{\bm{w}}_{i}$
@@ A.2 Multi-task parameter identification
We now come back to the multi-task setting, where the teacher generates a task-specific network using a hypernetwork conditioned on a task latent variable. Specifically, the weights of the second layer are fixed across the tasks, but the weights of the first layer are obtained as a linear combination of the modules: ${\bm{W}}(\Theta,{\bm{z}})=\sum_{m}^{M}z^{(m)}\Theta^{(m)}$. See section 3.1 for the full description.
@@
We provide results for MLPs with ReLU activation function as well as MLPs where the activation function belongs to a general class of smooth functions, for student networks that can be wider than the teacher network. Theorem 2 in the main text is a special case of the following results.
A.2.1 ReLU MLP
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $M\leq n$, $\hat{M}=M$, $-^\hat{h}\geq h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one ${\bm{z}}$ such that no two rows of $W(\Theta,{\bm{z}})$ are colinear. Finally, assume $\psi$ is the ReLU nonlinearity. Then, $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$ for almost any ${\bm{z}}$ is equivalent to the existence of an invertible matrix ${\bm{F}}$, a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ of $[1..\hat{h}]$ such that we have
@@
Assume there exists one point ${\bm{z}}_{0}$ such that $W(\Theta,{\bm{z}}_{0})$ has no two rows that are colinear. Then, for any linear subspace $U$ containing ${\bm{z}}_{0}$, $W(\Theta,{\bm{z}})$ has no two rows that are colinear for ${\bm{z}}\in U$ almost everywhere [^2] By ”${\bm{z}}\in U$ almost everywhere” we imply using the Lebesgue measure. In particular, the set of points ${\bm{z}}$ in $U$ such that $W(\Theta,{\bm{z}})$ has at least two colinear rows form a finite union of strict linear subspace of $U$.
@@
Recall that the $i$th row of $W(\Theta,z_{0})$ is given by $\Theta_{i}z_{0}$. Consider a pair $i,j$ with $i\neq j$. As no two rows in $W(\Theta,z_{0})$ are colinear, there exists only a finite number of $\alpha\in\mathbb{R}$ such that the equation $(\Theta_{i}-\alpha\Theta_{j}){\bm{z}}=0$ yields solutions of ${\bm{z}}$ that are not in Ker $\text{Ker }\Theta_{i}\cap\text{Ker }\Theta_{j}$. Given each of such $\alpha$, the set of ${\bm{z}}$ that verifies the equation form a linear subspace. Taking everything together, the set of ${\bm{z}}$ such that $W(\Theta,{\bm{z}})$ has at least two colinear elements is a finite union of linear subspaces. Due to the existence of ${\bm{z}}_{0}$ we know that these linear subspaces must not include any linear subspace containing ${\bm{z}}_{0}$ and the claim follows. ∎
@@
By assumption, there exists at least one point ${\bm{z}}_{0}$ such that $W(\Theta,{\bm{z}}_{0})$ has no two rows that are colinear. Following Lemma 6, for almost every ${\bm{z}}$ in $\mathbb{R}^{M}$, the first layer weights have no two rows that are colinear. For all ${\bm{z}}$, we fix $\hat{\bm{z}}\in\arg\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})$. By assumption, $L(\hat{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0$ for almost every ${\bm{z}}$. Considering the intersection of such ${\bm{z}}$ and those that generate first layer weights with no two colinear rows, we can apply Theorem 4 for almost every ${\bm{z}}$. For each such ${\bm{z}}$ the theorem guarantees us a partition ${\cal S}=({\cal S}_{1}(i),{\cal S}_{2}(i))_{i\in[1..h]}\oplus({\cal S}^{\prime}_{1}(i),{\cal S}^{\prime}_{2}(i))_{i\in[1..h^{\prime}]}\oplus({\cal S}_{0})$ such that equation 15-equation 22 hold for the generated weights.
@@
Because there exists a full rank matrix in the linear span of the $(\Theta_{i})_{1\leq i\leq h}$, then clearly we can find some scalar coefficients $(\lambda_{i})_{i\in[1..h]}$ and linearly combine the equations 46 over $i$ such that we obtain an equation of the form ${\bm{A}}{\bm{z}}={\bm{B}}\hat{\bm{z}}$ for all ${\bm{z}}\in\mathbb{R}^{M}$ and corresponding $\hat{\bm{z}}$, where ${\bm{A}}$ is full rank. Necessarily then, ${\bm{B}}$ is full rank, and multiplying by its pseudoinverse yields ${\bm{F}}{\bm{z}}=\hat{\bm{z}}$ where ${\bm{F}}$ is an invertible matrix. $\hat{\bm{z}}$ for which the equations hold is thus a linear function of ${\bm{z}}$, which also implies the constancy of the various parameters $\alpha$ w.r.t. ${\bm{z}}$.
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $M\leq n$, $\hat{M}=M$, $-^\hat{h}=h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one ${\bm{z}}$ such that no two rows of $W(\Theta,{\bm{z}})$ are colinear. Finally, assume $\psi$ is the ReLU nonlinearity. Then, $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$ for almost any ${\bm{z}}$ is equivalent to the existence of an invertible matrix ${\bm{F}}$, a partition ${\cal S}=({\cal S}_{1},{\cal S}_{2})$ and a permutation $\sigma$ of $[1..\hat{h}]$ such that we have
@@
A.2.2 MLP with infinitely differentiable activation function
@@
Assume a teacher and student two-layer network parameterized by resp. $({\bm{W}},{\bm{a}})$ and $(\hat{{\bm{W}}},\hat{{\bm{a}}})$, of hidden dimension resp. $-^h,\hat{h}$ with $-^h\leq\hat{h}$ and single output unit such that
@@
    • no output weight $a_{k}$ is 00
    • no input weight ${\bm{w}}_{k}$ is 00
@@
[…]
@@
In essence, if the teacher parameters are irreducible, we learn the same parameters up to permutation when the hidden layer is of the same width. If we allow the student to have a wider width, the overparameterization can either:
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $M\leq n$, $\hat{M}=M$, $-^\hat{h}\geq h$, and that there exists a full rank matrix in the linear span of $(\Theta_{i})_{1\leq i\leq h}$, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that there exists at least one ${\bm{z}}$ such that no two rows of $W(\Theta,{\bm{z}})$ are identical. Finally, assume the activation function $\psi$ satisfies $\psi\in\mathcal{C}^{\infty}$, and $\psi^{(n)}(0)\neq 0$ for infinitely many even and odd values of $n$. Then, $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$ for all ${\bm{z}}$ is equivalent to the existence of an invertible matrix ${\bm{F}}$ and a partition $(\mathcal{S}_{i})_{1\leq i\leq h}\oplus(\mathcal{S}^{\prime}_{i})_{1\leq i\leq h^{\prime}}$ for some $h^{\prime}$ of $-^[1,...\hat{h}]$ such that
@@
A.3 Example failure cases to build intuition
@@
We wish to uncover in the following the conditions under which a student might achieve compositional generalization when trained on a training set of tasks. This subsection provides some failure cases where perfectly learning the training tasks does not transfer to unseen tasks. It does not provide proofs, but rather an intuition on the later stated theorems and the assumptions they require.
In what follows, $(e_{m})$ denotes the canonical basis of $\mathbb{R}^{M}$.
@@
$\begin{array}[]{cccc}&m=1&m=2&m=3\\ \hat{\Theta}_{1}&A&B&0\\ \hat{\Theta}_{2}&0&B&C\\ \hat{\Theta}_{3}&0&B&0\\ \end{array}$
$\begin{array}[]{cc}&\text{output weight}\\ \hat{{\bm{w}}}^{(1)}&\lambda\\ \hat{{\bm{w}}}^{(2)}&\lambda\\ \hat{{\bm{w}}}^{(3)}&-\lambda\\ \end{array}$
@@
We can consider a student with two neurons and two modules. Let $\Theta_{1}=(A|B)$ and $\Theta_{2}=(C|D)$ for four linearly independent vectors $A,B,C,D$, and output weights all equal $a_{1}=a_{2}=\lambda$. If we only consider tasks defined by masks $\{e_{1},e_{2}\}$, a valid student is given by $\hat{\Theta}_{1}=(A|D)$ and $\hat{\Theta}_{2}=(C|B)$ and the same output weights. This student does however not generalize to the task described by mask $e_{1}+e_{2}$. Also see Figure A1 for a more detailed example of this failure case.
@@ A.4 Compositional generalization
We are now ready to state and prove the theorem. The proof for the general class of smooth activation function used in Appendix A.2 follows the same steps, we omit the details.
@@
Assume $\mathcal{P}_{x}$ has full support in the input space, $M\leq n$, $\hat{M}=M$, $-^\hat{h}=h$, that $\sum_{i}a_{i}\Theta_{i}$ is full rank, and $(a_{i})_{1\leq i\leq h}$ are all non-zero. Assume furthermore that for all modules $\Theta^{(m)}$, no two rows are colinear, that we can extract a family $Z_{k}$ from the support $\mathcal{Z}$ of $\mathcal{P}_{z}$ with the above property, and that $\mathcal{P}_{z}$ has connected and compositional support. Then,
$\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})\right]=0\implies\forall{\bm{z}},\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$ (61)
@@
$\forall{\bm{z}}\in\mathcal{F},\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{{\bm{z}}})=0\text{ }\implies\forall{\bm{z}},\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$ (62)
@@
[…]
@@
We start by arguing that we may assume without loss of generality that all $z\in Z_{k}$ satisfy some nice properties. First we claim that we may assume that all points $z\in Z_{k}$ satisfy $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$. This holds as $\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})\right]=0$ and our assumption that $\mathbb{P}_{{\bm{z}}\sim\mathcal{P}_{z}}[{\bm{z}}\in Z_{k}]>0$ implies that the set of points ${\bm{z}}$ of $Z_{k}$ such that $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})\neq 0$ has zero probability.
@@
We remind the reader that if a surjective function $f:A\rightarrow B$ between two metric spaces $A,B$ equipped by metrics $\mu,\mu^{\prime}$ verifies that all pre-image of nullsets are nullsets, i.e. $\forall X\subset B:\mu^{\prime}(X)=0\implies\mu(f^{-1}(X))=0$, then we have that if some property is true for almost every ${\bm{x}}\in B$, it must be true for $f({\bm{x}})$ for almost every ${\bm{x}}\in f^{-1}(B)=A$. If the above 2 points are true, then by application of this reasoning, we get that for almost any $({\bm{z}}_{e})_{e}\in S^{d+1}\setminus C$, there exists $(\lambda_{e})_{e}$ such that property 74 holds for $(\lambda_{e}{\bm{z}}_{e})_{e\in[1..d+1]}$. Since it can be shown that the measure of $S^{d+1}\setminus C$ is 00, we have that the assertion holds almost everywhere on $S^{d+1}$, which would conclude the proof.
@@
Let some $i=(i_{e})_{e}\in[1..h]^{d+1}$ and $s=(s_{e})_{e}\in\{-1,1\}^{d+1}$, such that there exists a pair $(n,m)$ such that $i_{n}\neq i_{m}$ or $s_{n}\neq s_{m}$. We fix such a pair. Let $M_{i,s}=(s_{1}\Theta_{i_{1}},..s_{d+1}\Theta_{i_{d+1}})$. We will now show that $\operatorname{Ker}{\cal I}\not\subset\operatorname{Ker}M_{i,s}$, i.e. that there exists a point in $\operatorname{Ker}{\cal I}$ which is not in $\operatorname{Ker}M_{i,s}$. Let the family $({\bm{z}}_{e})_{e}=(\mathbbm{1}_{e=n\lor e=m}(-1)^{\mathbbm{1}_{e=n}}{\bm{z}}_{0})_{e}$. Clearly, $\sum_{e}{\bm{z}}_{e}={\bm{z}}_{0}-{\bm{z}}_{0}=0$, thus $\Phi(({\bm{z}}_{e})_{e})$ belongs to the Kernel of ${\cal I}$. However, $\sum_{e}s_{e}\Theta_{i_{e}}{\bm{z}}_{e}=0\implies s_{n}\Theta_{i_{n}}{\bm{z}}_{0}=-s_{m}\Theta_{i_{m}}{\bm{z}}_{0}$. Because by assumption no two elements of $(\Theta_{i}{\bm{z}}_{0})_{i}$ are colinear, necessarily $\sum_{e}s_{e}\Theta_{i_{e}}{\bm{z}}_{e}=0\implies(s_{n}=s_{m})\land(i_{n}=i_{m})$, which contradicts the definition of $i,s$. Therefore, necessarily $\sum_{e}s_{e}\Theta_{i_{e}}{\bm{z}}_{e}\neq 0$ , i.e. $\Phi(({\bm{z}}_{e})_{e})\not\in\operatorname{Ker}M_{i,s}$.
@@
Fix some set $Z_{k}$. As in the continuous case we may assume without loss of generality that all points ${\bm{z}}$ in $Z_{k}$ satisfy that no two rows of ${\bm{W}}(\Theta,{\bm{z}})$ are colinear and that they achieve zero loss, i.e. $\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0$.
@@
[…]